# A star is born: Explosive Crump-Mode-Jagers branching processes in random environment

### v1

@@ Abstract.
We study a family of Crump-Mode-Jagers branching processes in random environment that explode, i.e. that grow infinitely large in finite time with positive probability. Building on recent work of the author and Iyer [10], we weaken certain assumptions required to prove that the branching process, at the time of explosion, contains a (unique) individual with an infinite offspring. We then apply these results to super-linear preferential attachment models. In particular, we fill gaps in some of the cases analysed in Appendix A of [10] and a study large range of previously unattainable cases.
Keywords: Explosive Crump-Mode-Jagers branching processes, super-linear preferential attachment, preferential attachment with fitness, random recursive trees, condensation.
@@ 1. Introduction
In a Crump-Mode-Jagers (CMJ) branching process (named after [4, 11]) an ancestral root individual produces offspring according to a collection of points on the non-negative real line. Each individual ‘born’ produces offspring according to an identically distributed collection of points, translated by their birth time. One is generally interested in properties of the population as a function of time. Classical work from the 1970s and ’80s related to this model generally deals with the Malthusian case, which, informally, refers to the fact that the population grows exponentially in time; see e.g. [3, 12] and the references therein for an overview of classical CMJ theory.
Far fewer results exist for CMJ branching processes when a Malthusian parameter does not exist. One family of CMJ branching processes for which this is the case, is the family of eplosive CMJ branching processes. Here, the total progeny of the branching process grows infinitely large in finite time with positive probability. Early work on such CMJ branching processes by Sevast’yanov [18, 19], Grey [5], Grishechkin [6], and Vatutin [20] studies necessary and sufficient conditions for (non-)explosion of Bellman-Harris processes (a special type CMJ branching processes). More recent work on this topic distinguishes between processes where individuals produce finitely many offspring in finite time almost surely, and processes where individuals can produce an infinite offspring themselves in finite time. In the former case, Komjáthy provides general criteria for explosion of CMJ branching processes in terms of solutions of a functional fixed point equation [13], and also extends the necessary and sufficient criteria for explosion in branching random walks in [1]. In the latter case, Oliveira and Spencer, and very recently Iyer [9], study a particular family of explosive CMJ branching processes with exponential inter-birth times [15], and Sagitov considers Bellman-Harris processes where individuals, upon death, may produce an infinite offspring [17, 16].
Recent work of the author and Iyer study a family of explosive CMJ branching processes in random environment [10]. Here, each individual born in the branching process is assigned an i.i.d. random weight, and the distribution of the point process that governs the offspring of the individual depends on their weight. Sufficient conditions for the emergence of infinite stars (individuals that produce an infinite offspring) and infinite paths (an infinite lineage of descendants) in the branching process, stopped at the (finite) time of explosion are formulated in a general set-up. Other sufficient criteria for explosion for CMJ processes in random environment are also studied in [8].
@@
Model definition. We consider individuals in the process as being labelled by elements of the infinite Ulam-Harris tree $\mathcal{U}_{\infty}:=\bigcup_{n\geq 0}\mathbb{N}^{n}$; where $\mathbb{N}^{0}:=\{\varnothing\}$ contains a single element which we call the root. We denote elements $u\in\mathcal{U}_{\infty}$ as a tuple, so that, if $u=(u_{1},\ldots,u_{k})\in\mathbb{N}^{k},k\geq 1$, we write $u=u_{1}\cdots u_{k}$. An individual $u=u_{1}u_{2}\cdots u_{k}$ is to be interpreted recursively as the $u_{k}^{\mathrm{th}}$ child of the individual $u_{1}\cdots u_{k-1}$; for example, $1,2,\ldots$ represent the offspring of $\varnothing$. Suppose that $(\Omega,\Sigma,\mathbb{P})$ is a complete probability space and $(S,\mathcal{S})$ is a measure space. We also equip $\mathcal{U}_{\infty}$ with the sigma algebra generated by singleton sets. Then, we fix random mappings $X:\Omega\times\mathcal{U}_{\infty}\rightarrow[0,\infty]$, $W:\Omega\times\mathcal{U}_{\infty}\rightarrow S$, and define $(X,W):\Omega\times\mathcal{U}_{\infty}\rightarrow[0,\infty]\times S$, so that $u\mapsto(X(u),W_{u})$. In general, for $u\in\mathcal{U}_{\infty}$ and $j\in\mathbb{N}$, one interprets $W_{u}$ as a ‘weight’ associated with $u$, and $X(uj)$ the waiting time or inter-birth times between the birth of the $(j-1)$th and $j$th child of $u$.
@@
Consequentially, a value of $X(ui)=\infty$ indicates that the individual $u$ has stopped producing offspring, and produces at most $i-1$ children.
We introduce some notation related to elements $u\in\mathcal{U}_{\infty}$. We use ||$\cdot$|| ⋅ | to measure the length of a tuple $u$, so that $|u|=0$ when $u=\varnothing$ and $|u|=k$ when $u=u_{1}\cdots u_{k}$. If, for some $x\in\mathcal{U}_{\infty}$, we have $x=uv$, we say $u$ is a ancestor of $x$. Further, given $\ell\leq|u|$, we write $u_{|_{\ell}}:=u_{1}\cdots u_{\ell}$ for the different ancestors of $u$. We equip $\mathcal{U}_{\infty}$ with the lexicographic total order $\leq_{L}$. Given elements $u,v$, we say $u\leq_{L}v$ if either $u$ is an ancestor of $v$, or $u_{\ell}<v_{\ell}$ where $\ell=\min\left\{i\in\mathbb{N}:u_{i}\neq v_{i}\right\}$. For $u\in\mathcal{U}_{\infty}$, we let $\mathcal{P}_{i}(u)$ denote the time, after the birth of $u$, required for $u$ to produce $i$ offspring. That is,
@@
For $t\geq 0$ we set $\mathscr{T}_{t}:=\{x\in\mathcal{U}_{\infty}:\mathcal{B}(x)\leq t\}$ and identify $\mathscr{T}_{t}$ as the genealogical tree of individuals with birth time at most $t$. For a given choice of $X,W$, we say $(\mathscr{T}_{t})_{t\geq 0}$ is the genealogical tree process associated with an $(X,W)$-Crump-Mode-Jagers branching process; often, we refer to $(\mathscr{T}_{t})_{t\geq 0}$ directly as an $(X,W)$-Crump-Mode-Jagers branching process, viewed as a stochastic process in $t$.
@@
where we adopt the convention that the infimum of the empty set is $\infty$. One readily verifies that $(|\mathscr{T}_{t}|)_{t\geq 0}$ is right-continuous, and thus $|\mathscr{T}_{\tau_{k}}|\geq k$. For each $k\in\mathbb{N}$ we define the tree $\mathcal{T}_{k}$ as the tree consisting of the first $k$ individuals in $\mathscr{T}_{\tau_{k}}$ ordered by birth time, breaking ties lexicographically. We call
@@
With the more commonly used notation for CMJ branching processes, one assigns a point process (denoted $\xi^{(u)}$) to each $u\in\mathcal{U}_{\infty}$, and refers to the points $\sigma^{(u)}_{1}\leq\sigma^{(u)}_{2},\ldots$ associated with this point process (in the notation used here $\mathcal{B}(u1),\mathcal{B}(u2),\ldots$). We do not use this framework here, because this requires one to be able to write the measure $\xi^{(u)}=\sum_{i=1}^{\infty}\delta_{\sigma_{i}}$, which requires one to impose $\sigma$-finiteness assumptions on the point process (see, for example, [14, Corollary 6.5]). This $\sigma$-finiteness is implied by the classical Malthusian condition, but, in this general setting, we believe it is easier to have a framework where one can directly refer to the points $\mathcal{B}(u1),\mathcal{B}(u2),\ldots$
Notation. Throughout the paper we use the following notation: We let :=$\mathbb{N}:=\{1,2,\ldots\}$ and set $\mathbb{N}_{0}:=\{0,1,\ldots\}$. For $x\in\mathbb{R}$, we let $\lceil x\rceil:=\inf\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor:=\sup\{n\in\mathbb{Z}:n\leq x\}$. For sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ such that $b_{n}$ is positive for all $n$, we say that $a_{n}=o(b_{n})$ and $a_{n}=\mathcal{O}(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$ and and if there exists a constant $C>0$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$, respectively. For random variables $X,(X_{n})_{n\in\mathbb{N}}$ we let $X_{n}\buildrel d\over{\longrightarrow}X,X_{n}\buildrel\mathbb{P}\over{\longrightarrow}X$ and $X_{n}\buildrel a.s.\over{\longrightarrow}X$ denote convergence in distribution, probability and almost sure convergence of $X_{n}$ to $X$, respectively. For a non-negative, real-valued random variable $X$ and $\lambda>0$, we let
@@
If the random variable $X$ has some dependence on a random variable $W\in S$, we write
$\mathcal{M}_{\lambda}(X;W):=\mathbb{E}\left[\exp(\lambda X)\,|\,W\right]\quad\text{and}\quad\mathcal{L}_{\lambda}(X;W):=\mathbb{E}\left[\exp(-\lambda X)\,|\,W\right].$
@@ 2. Results
Throughout this paper, we make the following general assumptions, that are common when studying (explosive) $(X,W)$-CMJ processes. We assume that the sequences of random variables $((X(uj))_{j\in\mathbb{N}},W_{u})$ are i.i.d. for different $u\in\mathcal{U}_{\infty}$. For a given $w\in S$, we let $(X_{w}(uj))_{j\in\mathbb{N}}$ denote a sequence $(X(uj))_{j\in\mathbb{N}}$, conditionally on the weight $W_{u}=w$. We also assume throughout that, for any $w\in S$ and any $u\in\mathcal{U}_{\infty}$, the random variables $(X_{w}(ui))_{i\in\mathbb{N}}$ are mutually independent.
@@ 2.1. Sufficient criteria for an infinite star in CMJ branching processes in random environment
    There exist positive random variables $(Y_{n})_{n\in\mathbb{N}_{0}}$ with finite mean such that, for any $w\in S$,
@@
    $\sum_{n=1}^{\infty}\mathcal{M}_{\lambda_{n}}(Y_{n})\mathbb{E}\left[\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n}(\varnothing);W)\right]<\infty.$ (2.2)
@@
    We have $\mathbb{E}\left[\sup\{k:X(k)=0\}\right]<1$. Additionally, for each $n\in\mathbb{N}$, almost surely
@@
In Assumption 2.1 we can consider Condition (a) as a uniform explosivity condition: regardless of its weight, the distribution of the time until an individual produces an infinite offspring, after having already produced $n$ children, is dominated by $Y_{n}$. This assumption is essential to work around dependencies that arise due to the random environment (i.e. the weights of individuals).
Condition (b) is used in the proof of Proposition 3.2, and provides an upper bound for the expected number of children that produce an infinite offspring before their parent does.
@@
$\limsup_{n\to\infty}\mathcal{M}_{\lambda_{n}}(Y_{n})<\infty\qquad\text{and}\qquad\ \sum_{n=1}^{\infty}\mathbb{E}\left[\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n}(\varnothing);W)\right]<\infty.$ (2.4)
@@
Under Assumption 2.1, the infinite tree $\mathcal{T}_{\infty}$ almost surely contains a node of infinite degree ((((an infinite star)))).
@@ 2.2. Application
    $\frac{f(\operatorname{deg}^{+}(j,T),W_{j})}{\sum_{i=1}^{n-1}f(\operatorname{deg}^{+}(i,T),W_{i})}.$ (2.5)
@@
[…]
@@
Let $(T_{i})_{i\in\mathbb{N}}$ be a preferential attachment tree with fitness with fitness function $f$, that satisfies Assumption ($w^{*}$). Let $(\lambda_{n})_{n\in\mathbb{N}}\subset[0,\infty)$ be increasing and tend to infinity with $n$, such that there exists $N\in\mathbb{N}$ with $\lambda_{n}<f(i,w^{*})$ for all $i\geq n\geq N$. If

### v2

# A star is born: Explosive Crump–Mode–Jagers branching processes

@@ Abstract.
We study a family of Crump–Mode–Jagers branching processes in random environment that explode, i.e. that grow infinitely large in finite time with positive probability. Building on recent work of the author and Iyer (“On the structure of genealogical trees associated with explosive Crump–Mode–Jagers branching processes”, arXiv:2311.14664, 2023), we weaken certain assumptions required to prove that the branching process, at the time of explosion, contains a (unique) individual with infinite offspring. We then apply these results to super-linear preferential attachment models. In particular, we fill gaps in some of the cases analysed in Appendix A of the work of the author and Iyer and study a large range of previously unattainable cases.
Keywords: Explosive Crump–Mode–Jagers branching processes, super-linear preferential attachment, preferential attachment with fitness, random recursive trees, condensation.
@@ 1. Introduction
In a Crump–Mode–Jagers (CMJ) branching process (named after [4, 11]) an ancestral root individual produces offspring according to a collection of points on the non-negative real line. Each individual ‘born’ produces offspring according to an identically distributed collection of points, translated by their birth time. One is generally interested in properties of the population as a function of time. Classical work from the 1970s and ’80s related to this model generally deals with the Malthusian case, which, informally, refers to the fact that the population grows exponentially in time; see e.g. [3, 12] and the references therein for an overview of classical CMJ theory.
Far fewer results exist for CMJ branching processes when a Malthusian parameter does not exist. One family of CMJ branching processes for which this is the case, is the family of explosive CMJ branching processes. Here, the total progeny of the branching process grows infinitely large in finite time with positive probability. Early work on such CMJ branching processes by Sevast’yanov [18, 19], Grey [5], Grishechkin [6], and Vatutin [20] studies necessary and sufficient conditions for (non-)explosion of Bellman-Harris processes (a special type CMJ branching processes). More recent work on this topic distinguishes between processes where individuals produce finitely many offspring in finite time almost surely, and processes where individuals can produce infinite offspring themselves in finite time. In the former case, Komjáthy provides general criteria for explosion of CMJ branching processes in terms of solutions of a functional fixed point equation [13], and also extends the necessary and sufficient criteria for explosion in branching random walks in [1]. In the latter case, Oliveira and Spencer, and very recently Iyer [9], study a particular family of explosive CMJ branching processes with exponential inter-birth times [15], and Sagitov considers Bellman-Harris processes where individuals, upon death, may produce infinite offspring [17, 16].
Recent work of the author and Iyer investigates a family of explosive CMJ branching processes in random environment [10]. Here, each individual born in the branching process is assigned an i.i.d. random weight, and the distribution of the point process that governs the offspring of the individual depends on their weight. Sufficient conditions for the emergence of infinite stars (individuals that produce infinite offspring) and infinite paths (an infinite lineage of descendants) in the branching process, stopped at the (finite) time of explosion are formulated in a general set-up. Other sufficient criteria for explosion for CMJ processes in random environment are also studied in [8].
@@
Model definition. We consider individuals in the process as being labelled by elements of the infinite Ulam-Harris tree $\mathcal{U}_{\infty}:=\bigcup_{n\geq 0}\mathbb{N}^{n}$; where $\mathbb{N}^{0}:=\{\varnothing\}$ contains a single element which we call the root. We denote elements $u\in\mathcal{U}_{\infty}$ as tuples, so that, if $u=(u_{1},\ldots,u_{k})\in\mathbb{N}^{k},k\geq 1$, we write $u=u_{1}\cdots u_{k}$. An individual $u=u_{1}u_{2}\cdots u_{k}$ is to be interpreted recursively as the $u_{k}^{\mathrm{th}}$ child of the individual $u_{1}\cdots u_{k-1}$; for example, $1,2,\ldots$ represent the offspring of $\varnothing$. Further, for individuals $u=(u_{1},\ldots,u_{k})$ and $v=(v_{1},\ldots,v_{\ell})$, we let $uv=(u_{1},\ldots,u_{k},v_{1},\ldots,v_{\ell})$ denote their concatenation. Suppose that $(\Omega,\Sigma,\mathbb{P})$ is a complete probability space and $(S,\mathcal{S})$ is a measure space. We also equip $\mathcal{U}_{\infty}$ with the sigma algebra generated by singleton sets. Then, we fix random mappings $X:\Omega\times\mathcal{U}_{\infty}\rightarrow[0,\infty]$, $W:\Omega\times\mathcal{U}_{\infty}\rightarrow S$, and define $(X,W):\Omega\times\mathcal{U}_{\infty}\rightarrow[0,\infty]\times S$, so that $(\omega,u)\mapsto((X(u))(\omega),W_{u}(\omega))$. In general, for $u\in\mathcal{U}_{\infty}$ and $j\in\mathbb{N}$, one interprets $W_{u}$ as a ‘weight’ associated with $u$, and $X(uj)$ the waiting time or inter-birth times between the birth of the (j1)(j-1)th and $j$th child of $u$.
@@
We note that a value of $X(ui)=\infty$ indicates that the individual $u$ has stopped producing offspring, and produces at most i1i-1 children.
We introduce some notation related to elements $u\in\mathcal{U}_{\infty}$. We use $|\cdot|$ to measure the length of a tuple $u$, so that $|u|=0$ when $u=\varnothing$ and $|u|=k$ when $u=u_{1}\cdots u_{k}$. If, for some $x\in\mathcal{U}_{\infty}$, we have $x=uv$, we say $u$ is a ancestor of $x$. Further, given $\ell\leq|u|$, we write $u_{|_{\ell}}:=u_{1}\cdots u_{\ell}$ for the different ancestors of $u$. We equip $\mathcal{U}_{\infty}$ with the lexicographic total order $\leq_{L}$. Given elements $u,v$, we say $u\leq_{L}v$ if either $u$ is an ancestor of $v$, or $u_{\ell}<v_{\ell}$ where $\ell=\min\left\{i\in\mathbb{N}:u_{i}\neq v_{i}\right\}$. For $u\in\mathcal{U}_{\infty}$, we let $\mathcal{P}_{i}(u)$ denote the time, after the birth of $u$, required for $u$ to produce $i$ offspring. That is,
@@
For $t\geq 0$ we set $\mathscr{T}_{t}:=\{x\in\mathcal{U}_{\infty}:\mathcal{B}(x)\leq t\}$ and identify $\mathscr{T}_{t}$ as the genealogical tree of individuals with birth time at most $t$. For a given choice of $X,W$, we say $(\mathscr{T}_{t})_{t\geq 0}$ is the genealogical tree process associated with an $(X,W)$-Crump–Mode–Jagers branching process; often, we refer to $(\mathscr{T}_{t})_{t\geq 0}$ directly as an $(X,W)$-Crump–Mode–Jagers branching process, viewed as a stochastic process in $t$.
@@
where we adopt the convention that the infimum of the empty set is $\infty$. One readily verifies that $(|\mathscr{T}_{t}|)_{t\geq 0}$ is right-continuous, and thus $|\mathscr{T}_{\tau_{k}}|\geq k$ (when $\tau_{k}<\infty$). For each $k\in\mathbb{N}$ we define the tree $\mathcal{T}_{k}$ as the tree consisting of the first $k$ individuals in $\mathscr{T}_{\tau_{k}}$ ordered by birth time, breaking ties lexicographically. We call
@@
With the more commonly used notation for CMJ branching processes, one assigns a point process (denoted $\xi^{(u)}$) to each $u\in\mathcal{U}_{\infty}$, and refers to the points $\sigma^{(u)}_{1}\leq\sigma^{(u)}_{2}$$\ldots$ associated with this point process (in the notation used here $\mathcal{B}(u1),\mathcal{B}(u2),\ldots$). We do not use this framework here, because this requires one to be able to write the measure $\xi^{(u)}=\sum_{i=1}^{\infty}\delta_{\sigma_{i}}$, which requires one to impose $\sigma$-finiteness assumptions on the point process (see, for example, [14, Corollary 6.5]). This $\sigma$-finiteness is implied by the classical Malthusian condition, but, in this general setting, we believe it is easier to have a framework where one can directly refer to the points $\mathcal{B}(u1),\mathcal{B}(u2),\ldots$
Notation. Throughout the paper we use the following notation: We let $\mathbb{N}:=\{1,2,\ldots\}$ and set $\mathbb{N}_{0}:=\{0,1,\ldots\}$ and $\mathbb{R}_{+}:=[0,\infty)$. For $n\in\mathbb{N}$, we set $n]:=\{1,\ldots,n\}$. For $x\in\mathbb{R}$, we let $\lceil x\rceil:=\inf\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor:=\sup\{n\in\mathbb{Z}:n\leq x\}$. For sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ such that $b_{n}$ is positive for all $n$, we say that $a_{n}=o(b_{n})$, $a_{n}=\mathcal{O}(b_{n})$, and $a_{n}=\Theta(b_{n})$ if $\lim_{n\to\infty}a_{n}/b_{n}=0$, if there exists a constant $C>0$ such that $|a_{n}|\leq Cb_{n}$ for all $n\in\mathbb{N}$, and if $a_{n}=\mathcal{O}(b_{n})$ and $b_{n}=\mathcal{O}(a_{n})$, respectively. For random variables $X,(X_{n})_{n\in\mathbb{N}}$ we let $X_{n}\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{d}}X,X_{n}\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{\mathbb{P}}}X$ and $X_{n}\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{a.s.}}X$ denote convergence in distribution, probability and almost sure convergence of $X_{n}$ to $X$, respectively. For a non-negative, real-valued random variable $X$ and $\lambda>0$, we let
@@ 2. Results
Throughout this paper, we make the following general assumptions, that are common when studying (explosive) $(X,W)$-CMJ processes. We assume that the pairs $((X(uj))_{j\in\mathbb{N}},W_{u})$ with $u\in\mathcal{U}_{\infty}$ are i.i.d. For a given $w\in S$, we let $(X_{w}(uj))_{j\in\mathbb{N}}$ denote a sequence $(X(uj))_{j\in\mathbb{N}}$, conditionally on the weight $W_{u}=w$. We also assume throughout that, for any $w\in S$ and any $u\in\mathcal{U}_{\infty}$, the random variables $(X_{w}(ui))_{i\in\mathbb{N}}$ are mutually independent.
@@ 2.1. Sufficient criteria for an infinite star in CMJ branching processes in random environment
    There exist positive random variables $(Y_{n})_{n\in\mathbb{N}_{0}}$ with finite mean such that, for any $w\in S$ and all $n\in\mathbb{N}_{0}$,
@@
    $\sum_{n=1}^{\infty}\mathcal{M}_{\lambda_{n}}(Y_{n})\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n}(\varnothing))<\infty.$ (2.2)
@@
    We have $\mathbb{E}\left[\sup\{k:X(1)=\cdots=X(k)=0\}\right]<1$. Additionally, for each $n\in\mathbb{N}$, almost surely
@@
In Assumption 2.1 we can consider Condition (a) as a uniform explosivity condition: regardless of its weight, the distribution of the time until an individual produces infinite offspring, after having already produced $n$ children, is dominated by $Y_{n}$. This assumption is essential to work around dependencies that arise due to the random environment (i.e. the weights of individuals).
Condition (b) is used in the proof of Proposition 3.2, and provides an upper bound for the expected number of children that produce infinite offspring before their parent does.
@@
$\limsup_{n\to\infty}\mathcal{M}_{\lambda_{n}}(Y_{n})<\infty\qquad\text{and}\qquad\ \sum_{n=1}^{\infty}\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n}(\varnothing))<\infty.$ (2.4)
@@
Under Assumption 2.1, the infinite tree $\mathcal{T}_{\infty}$ almost surely contains a node of infinite degree ((an infinite star)).
@@ 2.2. Application
    $\frac{f(\operatorname{deg}^{+}(j,T_{n-1}),W_{j})}{\sum_{i=1}^{n-1}f(\operatorname{deg}^{+}(i,T_{n-1}),W_{i})}.$ (2.5)
@@
$\{T_{i}:i\in\mathbb{N}\}\overset{d}{=}\{\mathcal{T}_{i}:i\in\mathbb{N}\}.$ (2.6)
Additionally, with $T_{\infty}:=\cup_{n=1}^{\infty}T_{n}$, we also have $T_{\infty}\overset{d}{=}\mathcal{T}_{\infty}$. As a result, the structural properties of $\mathcal{T}_{\infty}$ can be translated to $T_{\infty}$. The equality in distribution is a consequence of the memory-less property and the fact that the minimum of exponential random variables is also exponentially distributed, with a rate given by the sum of the rates of the corresponding variables; see for example [7, Section 2.1]. The use of continuous-time embeddings of combinatorial processes is pioneered by Athreya and Karlin [2].
@@
Let $(T_{i})_{i\in\mathbb{N}}$ be a preferential attachment tree with fitness function $f$ that satisfies Assumption ($w^{*}$). Let $(\lambda_{n})_{n\in\mathbb{N}}\subset[0,\infty)$ be increasing and tend to infinity with $n$, such that there exists $N\in\mathbb{N}$ with $\lambda_{n}<f(i,w^{*})$ for all $i\geq n\geq N$. If